when a fighter pilot makes a very quick turn, he experiences a centripetal acceleration. When this is greater than about 4 x g the pilot will usually lose consciousness out). Consider a pilot flying at a speed of 800 m/s ho wants to make a very sharp turn. What is the minimum radius of curvature that he can take without blacking out?
Answer #1

The min acceleration should be less than blacking out acceleration.

a = 4 g = 4 x 9.81 = 39.24 m/s^2

we know that, the centripital acceleration is given by:

a(c) = v^2/R => R = v^2/a

R = 800^2/39.24 = 16310 m

Hence, R = 16310 m